### 13 Apr 2026 v2

# 20XX Vol. X No. XX, 000–000

Abstract
We study outflow from an unmagnetized, shocked accretion disk around a non-rotating super-massive black hole using multidimensional hydrodynamics simulation with radiative cooling. We aim to investigate whether such shocked accretion flow can launch sustained collimated bipolar outflow reaching out to thousands of gravitational radii even in the absence of magnetic field and if yes, what terminal velocity can they achieve? We present the results of a few simulations of geometrically thick accretion flow with increasing specific angular momentum on a vertically elongated cylindrical domain. We show that bipolar outflow from a region very close to the black hole is originating and propagating vertically out to our simulation domain boundary at around $2651$ Schwarzschild radius. The outflow attains a terminal velocity with a maximum value found to be $0.14c$ and the outflow rate depends on the angular momentum value of the accreting material. We also compute the self-Comptonized bremsstrahlung spectra for all the disk-jet runs.
keywords:
accretion, accretion disks — black hole physics — hydrodynamics — shock waves — radiative transfer — methods: numerical
1 Introduction
Many high-energy sources containing black hole show wind and/or jet (Mirabel and Rodríguez 1994; Stirling et al. 2001; Gallo 2010; Blandford et al. 2019). Whether the black hole is part of a binary system or located at the core of a galaxy, the qualitative structure of outflow is very similar when it is present. The outflow originates from very close to the central object and extends from hundreds to thousands of gravitational radii. It is well understood that, unlike a normal star, a black hole itself does not eject matter. Rather, the gravitationally captured matter, forming an accretion disk around the black hole, produces the outflow. Given that the gravitational attraction is very strong close to a black hole, it is thus very surprising how a fraction of the infalling matter gets rid of this strong attraction and becomes unbounded.
Outflow formation from the magnetized plasma of an accretion disk is mainly attributed to two mechanisms: Blandford-Znayek (Blandford and Znajek 1977) and Blandford-Pyne (Blandford and Payne 1982). It is, of course not very surprising that the accreting material is magnetized. Magnetic field is very pervasive and wanders along with plasma. In a binary system, matter can drag in magnetic field from the companion star. In the case of accretion onto a galaxy center, the infalling matter can drag magnetic field from nearby stars or from the interstellar medium. As this magnetized matter moves closer to the black hole, the magnetic field structure, as well as the magnitude may change due to various physical processes such as geometric compression, dynamo effect, magnetic reconnection, etc. This magnetic field can then help in shaping the accretion disk structure and launching outflow from the accretion disk in the manner described in these references.
However, the magnetic field may not be the only driving force behind launching outflow. In principle, if a sufficiently strong force develops in a way that cancels the gravitational pull, matter can be driven away from the accretion disk. Thus, if the outward centrifugal force due to rotating matter becomes sufficiently strong, the effective potential barrier (Misner et al. 1973; Shapiro and Teukolsky 1983) may be so high that the matter may not be able to accrete onto the central black hole and be driven away. Also, a sufficiently strong pressure gradient force acting opposite to the gravitational attraction can drive material away. The pressure can be due to radiation or thermal. If the disk material and the radiation are interacting effectively, radiation pressure can act in that case (Icke 1989; Fukue 1996; Chattopadhyay et al. 2004; Vyas et al. 2015). On the other hand, if the cooling time scale is lower than the infall time scale, matter can become sufficiently hot and the required thermal pressure gradient force can develop inside the accretion flow (e.g., Blandford and Begelman 1999; Soker 2023). If an accretion disk, having satisfied such conditions, additionally carries magnetic field, all the effects may add up to make the outflow launching favorable (e.g., Fukue 2004).
Outflows are not observed in isolation. Generally, spectral states are correlated to the outflow activities for X-ray binaries as well as active galactic nuclei (AGNs). Outflow is observed to be enhanced during harder spectral states (Fender et al. 2000; Done et al. 2007; Belloni 2010; Fender and Belloni 2012; Zdziarski et al. 2014). Thus, while focusing on the theoretical understanding of the outflow formation, one should pay attention to the fact that the same flow solution has to produce a consistent spectral state.
There have been several analytical and simulation studies of rotating, transonic accretion flow, which is sub-Keplerian far away from the black hole (Fukue 1987; Chakrabarti 1989; Chakrabarti and Molteni 1993; Ryu et al. 1995; Molteni et al. 1996a; Chakrabarti 1996; Das et al. 2014; Lee et al. 2016; Kim et al. 2017; Suková et al. 2017; Kim et al. 2019; Garain and Kim 2023; Debnath et al. 2024). Being sub-Keplerian, the gravity is dominating far away. Hence, the matter initially moves to the black hole with nearly free-falling radial velocity and becomes supersonic. However, for such a rotating accretion flow, the outward centrifugal force can become comparable to the inward gravitational pull closer to the black hole. As a result, the free fall of matter is slowed down at a certain radius (few tens to hundreds of Schwarzschild radii) and the matter starts piling up. If the centrifugal force is sufficiently strong, the flow can experience a shock. However, this mechanism differs from the boundary-layer shock model (Soker and Regev 2003), which applies to young stellar objects and involves shocks formed where the Keplerian disk interacts with the stellar surface. In contrast, our model concerns accretion onto black holes, where no physical surface exists, and the shocks arise dynamically from angular-momentum-induced centrifugal effects within the flow. Such an accretion solution is shown to produce outflow and outflow rate dependence on various inflow parameters are studied analytically and numerically (Molteni et al. 1994; Ryu et al. 1997; Chakrabarti 1999; Singh and Chakrabarti 2011; Garain et al. 2012; Aktar et al. 2015). The postshock region becomes hotter, denser and geometrically thick. Bipolar outflow can originate from this region due to the combined effects of centrifugal and pressure gradient forces. If the accreting matter additionally brings in magnetic field, it becomes stronger in the post-shock region due to compression and contribute positively to the outflow launching (Deb et al. 2017; Mondal and Mukhopadhyay 2019; Shende et al. 2019; Garain et al. 2020). The postshock flow becomes the base of the outflow.
This disk-jet solution simultaneously addresses the spectro-temporal states. Increased density inside this postshock flow makes this region optically thicker. As a result, it can intercept photons originating within the accretion disk and Comptonize or inverse-Comptonize them. In the popularly known two-component advective flow (TCAF, Chakrabarti and Titarchuk 1995; Mondal and Chakrabarti 2021) model, such sub-Keplerian flow forms one component, and a Keplerian disk component is sandwiched inside this. Lower energy seed photons from the Keplerian disk is inverse-Comptonized to higher energy by this post-shock region and produce harder spectrum. During this time, outflow is also launched from the post-shock region (Garain et al. 2012, 2014). However, when the sub-Keplerian component is nearly absent, the seed photons leave unaffected, producing a softer state and outflow is also absent. In the case when the flow has single component (i.e., only sub-Keplerian matter), even then, the seed photons generated inside this component through the bremsstrahlung or synchrotron process can get self-Comptonized and produce harder radiations (Chakrabarti and Mandal 2006; Mandal and Chakrabarti 2008). In such situations, outflow may accompany the harder state.
In this paper, we aim to investigate the fate of the bipolar outflow launched from the postshock region, at a very large distance from the black hole. We do not consider any magnetic field just to focus on the effects of centrifugal and thermal pressure gradient forces on the outflow. We are specifically interested to see whether such a shocked accretion flow can launch sustained collimated outflow reaching out to thousands of gravitational radii and if yes, what terminal velocity can they achieve? We present the results of several simulations with increasing the specific angular momentum to explore the effect of increasing centrifugal force on the outflow base (i.e., postshock region) as well as on the outflow rate at far distance. We additionally generate seed bremsstrahlung photons inside the disk-jet system and compute the self-Comptonized spectra of these seed photons using a monte-carlo based radiative transfer simulation (Pozdnyakov et al. 1983b; Ghosh et al. 2009, 2010; Garain et al. 2014).
In this paper, we choose $r_{g}=2GM_{\rm{bh}}/c^{2}$ as the unit of distance, $c$ as unit of velocity, $r_{g}c$ as unit of angular momentum, and $r_{g}/c$ as unit of time, unless specified otherwise. Here, $G$ is the gravitational constant, $M_{\rm{bh}}$ is the mass of the black hole, and $c$ is the speed of light in free space.
2 Simulation procedure
For our present study, we solve the inviscid hydrodynamics equations and include thermal bremsstrahlung cooling in the energy conservation equation as a source term. We assume axisymmetry of the accretion disk. Thus, the system can be described in two-dimensional cylindrical coordinates $R,Z$ as follows:
$\frac{\partial\mathbf{U}}{\partial t}+\frac{1}{R}\frac{\partial\left(R\mathbf{F_{R}}\right)}{\partial R}+\frac{\partial\mathbf{F_{Z}}}{\partial Z}=\mathbf{S},$ (1)
where, the vector of conserved variables $\mathbf{U}$, R-flux $\mathbf{F_{R}}$, and Z-flux $\mathbf{F_{Z}}$ can be written as:
${\bf U}=\left(\begin{array}[]{ccccc}\rho\\ \rho v_{R}\\ \rho l\\ \rho v_{Z}\\ E\end{array}\right);\quad{\bf F_{R}}=\left(\begin{array}[]{ccccc}\rho v_{R}\\ \rho v_{R}^{2}+P\\ \rho lv_{R}\\ \rho v_{R}v_{Z}\\ \left(E+P\right)v_{R}\end{array}\right);\quad{\bf F_{Z}}=\left(\begin{array}[]{ccccc}\rho v_{Z}\\ \rho v_{R}v_{Z}\\ \rho lv_{Z}\\ \rho v_{Z}^{2}+P\\ \left(E+P\right)v_{Z}\end{array}\right)$ (17)
And the source term is:
${\bf S}$ $=$ $\left(\begin{array}[]{ccccc}0\\ \frac{\rho v_{\phi}^{2}}{R}+\frac{P}{R}-\rho\frac{\partial\Phi}{\partial r}\frac{R}{r}\\ 0\\ -\rho\frac{\partial\Phi}{\partial r}\frac{Z}{r}\\ -\rho\frac{\partial\Phi}{\partial r}\frac{\left(Rv_{R}+Zv_{Z}\right)}{r}-Q_{br}\end{array}\right).$ (23)
Here, $\rho$ is mass density, $v_{R},v_{\phi},v_{Z}$ are the $R$, $\phi$ and $Z$ components of velocity, $P$ is pressure, $E=\frac{1}{2}\rho(v_{R}^{2}+v_{\phi}^{2}+v_{Z}^{2})+\frac{P}{\gamma-1}$ and $l=Rv_{\phi}$. $r$ represents the spherical radius and is given by $r=\sqrt{R^{2}+Z^{2}}$. $\Phi$ represents the gravitational potential (Paczyńsky and Wiita (1980)): $\Phi=-GM_{\rm{bh}}/(r-r_{g})$. We use a polytropic equation of state $P=K\rho^{\gamma}$ with $\gamma=4/3$.
Inside the simulation domain, $\rho$ is measured with respect to the reference density $\rho_{0}$ at the outer radial boundary:
$\rho_{0}=\frac{\dot{m}}{2\pi R_{\text{out}}\,h\,v_{R0}}$ (24)
Here, $\dot{m}$, $h$ and $v_{R0}$ are the mass injection rate, thickness of the disk, and injection radial velocity, respectively, at $R=R_{\mathrm{o}ut}$. In the chosen unit system, bremsstrahlung loss $Q_{br}$ is given as (Molteni et al. (1996b))
$Q_{\mathrm{br}}=1.43\times 10^{-27}\,\rho^{2}\left(\frac{P}{\rho}\right)^{1/2}\,\frac{\rho_{0}r_{g}T_{\mathrm{ref}}^{1/2}}{m_{p}^{2}c^{3}}$ (25)
Here, $m_{p}$ is mass of a proton and $T_{\mathrm{ref}}$ is given by
$T_{\mathrm{ref}}=\frac{\mu m_{p}c^{2}}{k_{B}}$ (26)
$\mu$ = 0.5 and $k_{B}$ being the Boltzmann constant.
We solve the above equations using the code described in Garain and Kim (2023). We use the two-dimensional version of that code. It is a finite-volume method-based globally second-order accurate code. van-Leer slope limiter (Mignone 2014) has been used for spatial reconstruction, HLL Riemann solver has been used for interfacial flux calculation, and two-stage strong-stability preserving Runge-Kutta (Shu and Osher 1988) method is used for time integration. For details, please see the above references. The source terms are treated using an explicit method.
2.1 Simulation set-up
Our simulation set-up is very much similar to those which have been used over decades to study such shocked accretion flow (Chakrabarti and Molteni 1993; Ryu et al. 1995; Molteni et al. 1996a; Chakrabarti et al. 2004; Okuda et al. 2007; Giri et al. 2010; Okuda and Molteni 2012; Das et al. 2014; Lee et al. 2016; Kim et al. 2017, 2019; Garain and Kim 2023; Debnath et al. 2024). We perform our simulations on a vertically elongated RZR-Z domain $0:250]\times[-2651:2651]$. All the simulations are run using $480\times 2000$ ratioed grid points having a common ratio $1.0043$. Grids are constructed such that we resolve the region close to the origin with the highest resolution and grids become coarser as we move away. Thus, along the $R$ direction, finer grids are located close to $R=0$ and along the $Z$ direction, ratioing is done symmetrically about the equatorial plane. The black hole is placed at the origin, and we apply an absorbing boundary condition within a radius $r=1.8$. Mass of the black hole is chosen to be $_{\rm{bh}}=6.5\times 10^{9} M. M$ass injection rate $\dot{m}=0.01$ mass Eddington rate is assumed. We present the results of a total of five simulations here.
ID $l$ $v_{R0}$ $a_{R0}$ Terminal velocity
A1 1.50 0.03015 0.02907 0.05
A2 1.60 0.03006 0.02907 0.09
A3 1.70 0.02996 0.02907 0.12
A4 1.75 0.02991 0.02907 0.14
A5 1.80 0.02986 0.02908 0.1
Table 1: Simulation parameters where $l$, $v_{R0}$ and $a_{R0}$ represent the specific angular momentum of the accretion flow, radial velocity and sound speed of the injected matter respectively. Last column represents the terminal velocity of the outflow. See text for details.
Table 1 shows the simulation parameters presented in this paper. Column 1 shows the case IDs. Column 2 shows the specific angular momentum of the accretion flow $l$ at the outer boundary. Columns 3 and 4 show the radial velocity $v_{R0}$ and the sound speed $a_{R0}$ of the injected matter at the outer boundary. These are computed following the theoretical calculation provided in Chakrabarti (1989). The calculation requires two conserved parameters, specific energy $\epsilon=\frac{1}{2}(v_{R}^{2}+v_{\phi}^{2}+v_{Z}^{2})+\frac{a^{2}}{\gamma-1}+\Phi$ ($a=\frac{\gamma P}{\rho}$ being the local sound speed) and specific angular momentum $l$ of the flow, to figure out the radial variation of the flow profile. For all our runs, we choose $\epsilon=0.001$. This $\epsilon$, along with $l$ (Column 2), determines $v_{R0}$ and $a_{R0}$ at $R=R_{out}$. Our choice of ($l,\epsilon$) is guided by the parameter space classification of this flow, Chakrabarti and Das (2001) which shows that for these sets, shocks can develop in the accretion solution. Additionally, we wish to investigate the effect of increased centrifugal barrier on the outflow and hence increase $l$ keeping $\epsilon$ constant. The last column represents the terminal velocities attained by the outflowing matter in the respective simulation. All simulations are run until a time stop of 160000. The solutions have achieved steady states by this time.
2.2 Initial and boundary conditions
We start our simulation with the entire domain initially filled with static matter having constant floor density $\rho_{\mathrm{floor}}=10^{-6}$ and floor pressure $P_{\mathrm{floor}}=\frac{a_{R0}^{2}}{\gamma}\rho_{\mathrm{floor}}$. Accreting matter enters the simulation domain axisymmetrically through the outer radial boundary $R=R_{\mathrm{out}}=250$. At this inflow boundary, the disk is assumed to be thick, and the disk half-height is $h/2=280$. Therefore, the inflow boundary condition is applied for $h/2\leq Z\leq h/2$ at $R=R_{\mathrm{o}ut}$. At the ghost grids adjacent to all the active grids within this $Z$ range, we maintain time-independent matter density $\rho_{0}=1$, velocity $v_{R0}$ (see Table 1) and pressure $P_{0}=a_{R0}^{2}\rho_{0}/\gamma$ throughout the simulation duration. The idea behind this is that the mass inflow rate towards the black hole remains unchanged throughout the simulation duration. At other $Z$ at $R=R_{\mathrm{o}ut}$, we apply outflow boundary condition. Reflection boundary condition has been imposed on the axis $R=0$. At the upper and lower $Z$ boundaries, outflow boundary conditions are enforced.
2.3 Bremsstrahlung emission modeling and self-Comptonization
Inside an accretion disk around a super-massive black hole, thermal bremsstrahlung (free–free emission) constitutes an important radiative mechanism contributing to X-ray and $\gamma$-ray emissions. Since we perform simulations of geometrically thick accretion disks ($h(R)\sim R$, where $h$ is the accretion disk half-height at radius $R$) with small $\dot{m}$, we have low optical depth values (optical depth of 0.15 cm-1 as shown in Section 3.3) Additionally, majority($\approx 99\%$) of the grid points in our simulation domain have temperature $2\times 10^{7}K$, hence we consider bremsstrahlung to be the dominant radiative mechanism and ignore line emissions. The spectral emissivity (erg/cc/sec/Hz) due to bremsstrahlung emission is given by (Rybicki and Lightman 1991):
$\frac{dW}{dV\,dt\,d\nu}=6.8\times 10^{-38}\rho^{2}T^{-1/2}\exp\left(-\frac{h\nu}{k_{B}T}\right)$ (27)
By integrating this expression over frequency, the total thermal bremsstrahlung emissivity, as given in Equation (25), can be obtained. These expressions form the foundation for modeling thermal bremsstrahlung emission in high-temperature plasma in the vicinity of black holes.
To obtain the spatial distribution of photons, we use a Monte Carlo sampling technique based on the cumulative distribution function (CDF)(Sobol 1994; Ross 2009). The total emission power (erg/sec) from each grid cell is first normalized to construct a discrete probability distribution, where each element represents the probability of a photon being emitted from the corresponding cell. The cumulative sum of this probability distribution refers to the CDF, which partitions the unit interval [0, 1] into segments proportional to the respective emission probabilities. A fixed number of random numbers are then drawn between [0, 1]. Each number is then mapped to the grid point corresponding to the CDF interval in which it falls, i.e., a random number smaller than the CDF value of the first cell is assigned to the first grid cell, while values ranging between two successive CDF values are assigned to the corresponding grid cell. In this way, the fraction of photons associated with each grid cell statistically reproduces the underlying emission probability distribution, with the accuracy improving as the number of sampled photons increases. The weight factor for modeling is calculated by dividing the integrated emission from each cell by the number of Monte Carlo photons distributed in that cell.
The seed photons for Comptonization are generated using the weighted photon distribution described above, followed by modeling the energy distribution through Gamma function modeling (Pozdnyakov et al. 1983c; McGrath and Irving 1975) using a single random number.
Once a photon is generated within the disk-outflow configuration, following the above algorithm, we track the photon until it leaves the entire computational domain, or it is absorbed by the black hole. During its tracking, if it satisfies a scattering condition (i.e., its optical depth becomes larger than a threshold value), a (inverse-)Comptonization is modeled. For modeling these tracking and Comptonization processes, we follow the same methods used in Pozdnyakov et al. (1983a); Ghosh et al. (2009, 2010); Garain et al. (2012, 2014). We also consider gravitational red-shift during the tracking of the photons. The spectrum is calculated in a post-processing manner on the final disk-outflow configuration. Thus, we do not include the heating/cooling effects due to the Comptonization part in the time-dependent fluid dynamics. For our current model parameters, we observe that a very small fraction ($1\%$) is intercepted by the disk-outflow configuration. Thus, we believe this exclusion of heating/cooling effects won’t affect the fluid dynamics significantly. We use 108 seed photons for the spectral modeling. This number has been chosen based on a convergence test. We ensure that the final spectrum is converged and Monte Carlo noise is upto the level of tolerance.
3 Results
3.1 Flow dynamics
Figure 1: shows the mass density ($\rho$) distribution (in color) on log scale overlaid with velocity arrows at the final time for all cases A1-A5 (a:A1, b:A2, c:A3, d:A4, e:A5). Black color shows lower densities, and red color shows higher densities. Length of a velocity arrow is proportional to the magnitude $\sqrt{v_{R}^{2}+v_{Z}^{2}}$. Colors of arrows represent their magnitudes as shown in the legend for velocity vector. The maximum velocity is found extremely close to the black hole on the equatorial plane
Fig.  1 shows the density-velocity distribution at the final time for all the runs A1-A5 (a:A1, b:A2, c:A3, d:A4, e:A5). Colors represent the mass density on a log scale and the velocity field ($v_{R},v_{Z}$) are overplotted on this as arrows. Black color shows lower density ($\approx 10^{-5}$ w.r.t. reference density $\rho_{0}$) and red color shows higher density. Length of an arrow is proportional to the velocity magnitude $v_{r}=\sqrt{v_{R}^{2}+v_{Z}^{2}}$. Although the simulations are conducted inside the radial domain $0:R_{\mathrm{o}ut}$], we have reflected results about the vertical axis while plotting. Matter enters the simulation domain at $R_{\rm out}=250$ with the velocity vectors pointed towards the black hole located at the center. After the transient phase is over (by time $t\sim 20000$), the simulation domain is mostly filled with disk-outflow matter. We observe that the region near the equator is filled with a geometrically thick accretion disk, while regions vertically up and down are filled with outflowing matter. The outflow is hollow in structure, meaning the region along the axis is void of matter.
Figure 2: (a) shows a zoomed-in view of case A5 in the RZR-Z domain $0:150]\times[-160:160]$. We plot the number density (measured in particle number/$c$) colormap overlaid with velocity arrows. The iso-density contour (green line) corresponding to the value $8.85\times 10^{6}$, allows to identify the shock surface at $R\sim 100$. The region inside this contour is the high-density post-shock region, and the outer low-density region is the pre-shock region, marked accordingly. The post-shock region shows the presence of turbulent vortices. We also mark the jet launching region close to the black hole. (b) This figure shows the temperature (in Kelvin) distribution in the same domain clearly separating the high-temperature post-shock region from the low-temperature pre-shock region for case A5.
Zoomed-in observation of the equatorial region reveals the thick disk structure as well as the outflow launching mechanism. Fig  2 shows the central region $0:150]\times[-160:160]$ for the case A5. Fig  2(a) shows the number density distribution and (b) shows the temperature distribution. We notice that as the infalling matter moves towards the central region, density (and temperature as well) increases due to geometric compression initially and then due to a shock discontinuity. The discontinuity structure bends outward away from the equatorial plane (Molteni et al. 1994, 1996a) giving rise to a centrifugal pressure-supported boundary layer (CENBOL). Matter density increases past this discontinuity surface again due to geometric compression. An iso-density contour plot may allow one to see the geometrically thick, toroidal disk structure in this region. The green iso-density contour corresponding to number density value $8.85\times 10^{6}$ in Fig  2(a) shows such a toroidal structure. Velocity vectors in the post-shock region show presence of turbulent vortices as well. We also notice from the velocity vectors that outflow is primarily launched from this post-shock region. This observation holds for all the cases. However, the centrifugal force varies from one case to another and that affects the shock strength and, hence, the compression. This causes the results to differ. We provide further analysis below to quantify these.
Figure 3: represents the radial variations (along the equator) of logarithm of temperature($T$), number density ($n$), and velocity ($v_{R}$) at the final time. Different quantities are translated by a scalar along the vertical axis to bring them to the same scale. We show temperature in Kelvin, number density in particle number/cc and velocity in cm/s units.
The discontinuous jump and subsequent compression become prominent when we plot the radial variation of different fluid variables. Figure 3 shows the radial variation of number density, velocity, and temperature along the equatorial line for all the cases at the final time. Density and temperature are measured in particle number/cc and Kelvin units, respectively. For the low angular momentum cases ($\lambda=1.50$ and $1.60$, corresponding to runs A1 and A2 respectively), the flow shows a mild deceleration (velocity drop) at the centrifugal barrier around $R=28$ (A1) and $R=32$ (A2). A density compression (compression ratio $\rho_{\rm post-shock}/\rho_{\rm pre-shock}\sim 6.25$(A1) and $\sim 6.28$(A2)) and temperature enhancement is seen due to this. For slightly higher angular momentum $\lambda=1.70$ (run A3), the centrifugal barrier becomes strong enough to produce a well-defined shock (supersonic to subsonic transition) at $R=34.5$, resulting in a sharp, larger drop in velocity. This causes larger density compression and temperature rise. The density compression ratio is found to be $7.4$ in this case. We also observe the presence of an inner shock near $R=5$. For $\lambda=1.75$ (run A4), the shock feature is more pronounced (shock at $R=34.8$), and the density compression ratio is found to be $8.7$. The flow also exhibits evidence of a secondary, inner shock closer to the black hole ($R=13.5$) before passing through the inner sonic point. For the highest angular momentum $\lambda=1.80$ (run A5) case, the flow shows unsteady behavior, both radial and vertical oscillations of the post-shock matter are observed. Careful observation of successive snapshots show the formation and disappearance of turbulent vortices of different sizes in the post-shock matter. Since the centrifugal barrier is very strong in this case, the bounced-back matter from this barrier interacts with the incoming matter, resulting in the initial formation of large vortices, above and below the equator. Subsequent velocity shear between the the incoming flow and the matter in these vortices develop shear instability. This instability possibly grows in time, resulting in the formation of turbulent flow. Thus, the post-shock region becomes highly turbulent with many intermediate discontinuities. The radial profiles of velocity, number density, and temperature at the final time for this case show such discontinuities. Such unsteady shock dynamics for high angular momentum flow have been reported in earlier numerical simulations (Ryu et al. 1997; Deb et al. 2016; Garain and Kim 2023).
The CENBOL structure is nearly stationary for the low angular momentum runs (A1 and A2), while cases A3 and A4 exhibit quasi-periodic radial oscillations associated with shock transitions. For the turbulent, high angular momentum case (A5), the boundary layer is highly dynamic, showing significant temporal variability. In Fig. 4, we plot the time variations of the CENBOL at the equator for all five cases. Time is represented in years here. With increasing centrifugal force, the average boundary location moves farther from the black hole.
Figure 4: shows temporal variation of the CENBOL boundary at the equator for all cases. Cases A1 and A2 do not show shock formation, though a discontinuity due to centrifugal barrier is found. Cases A3, A4, and A5 show shock formation and oscillations of the CENBOL boundary over time.
3.2 Outflow
The dynamical CENBOL is found to launch outflow from a region close to the black hole. At the boundary layer, flow is slowed down, and as a result, the kinetic energy of the flow is converted to thermal energy. Thus, inside the CENBOL, the flow becomes hotter. Further geometric compression makes the rotating flow even hotter. Thus, the flow develops a pressure gradient in both radial and vertical directions. This vertical pressure gradient force launches outflow of matter in the vertical direction. Since the ram pressure of the infalling matter prevents matter from expanding in the radial direction, there is no outflow in the radial direction. Thus, we see only bipolar outflow being launched in all the cases.
To investigate whether the outflowing matter can actually become unbounded, we plot the specific energy($\epsilon$) distribution inside the disk-outflow region in Fig. 5(a). These plots have been done using the final timestep data. For all the runs, infalling matter has $\epsilon=10^{-3}$. This value is mostly conserved within the accretion disk of thickness 280<Z<280-280<Z<280 at all radii except very close to the black hole. Inside the compressed region, energy is redistributed because of the turbulent interaction between the incoming and the centrifugal barrier reflected matter. Thus, we observe some matter even with a negative $\epsilon$. Such matter will be ultimately absorbed by the black hole. Importantly, we observe that the outflowing matter generally has higher values of the $\epsilon$ and is always positive, making sure that this matter will contribute to the unbound outflow.
[…]
Figure 6: shows time-averaged outflow velocity as a function of radial distance from the origin for all the cases. Outflowing matter is found to attain a terminal velocity after it moves a certain distance from the black hole.
Fig.  6 shows the time-averaged vertical velocity ($v_{z}$) of the outflowing matter as a function of radial distance from the central black hole for different cases. The time averaging has been performed over the entire steady-state duration. From velocity vectors in Fig.  1, we note that within $|Z|\lesssim 300$, the outflowing region has a thin width. The velocity shear between outflowing and inflowing matter in this region develops Kelvin-Helmholtz instability. Thus, we do not get reliable values of $v_{z}$ in this region. As matter moves away beyond $|Z|\sim 300$, the outflow expands radially as the vertical disk thickness of the inflowing accreting matter is less than 300. The outflow is mostly hollow in structure, and the inner surface is highly dynamic. Thus, while calculating the $v_{z}$ at a given distance, we ensure that the distant point is well inside the outflow region.
[…]
We also quantify the outward mass flux through the top and the bottom surfaces ($|Z|=2651$) of the computational domain. Fig.  7 shows the total mass flux through these two surfaces, normalized by the mass injection through the outer radial boundary. On average, we find the highest mass outflow rate ($\sim$ 0.04 % of the injection rate) for the highest $l$ (run A5) case. We emphasize that the measured outflow rate is within $\sim$ 5 opening angle of the jet.
Figure 7: shows the outflow rate, normalized by the respective injection rate, over time.
3.3 Radiative properties
Figure 8: shows the bolometric bremsstrahlung luminosity as a function of time for all accretion cases A1–A5. All our simulations are done for a black hole of mass of $M=6.5\times 10^{9}\,M_{\odot}$ with a mass accretion rate of $\dot{m}=0.01$ (in units of the mass Eddington rate), corresponding to a physical mass accretion rate of $\dot{M}\approx 9.36\times 10^{24}\,\mathrm{g\,s^{-1}}$ ($\approx 0.15\,M_{\odot}\,\mathrm{yr^{-1}}$). The resulting bolometric luminosity is of the order of $10^{42}$ for the case A5   $\mathrm{erg\,s^{-1}}$, yielding a radiative efficiency of $L_{\mathrm{bol}}/(\dot{M}c^{2})\approx 4.29\times 10^{-4}$.
Figure 9: shows the net (seed plus the Comptonized) bremsstrahlung spectra for all the cases. See text for more details.
Fig.  8 shows the temporal evolution of the bolometric bremsstrahlung luminosity (erg/sec)for the five cases marked in the legend. Time is measured in years. The bremsstrahlung emission is computed as the volume-integrated emissivity of the plasma using Eq. 25 and primarily contributed by the density and temperature distribution inside the CENBOL region. For the lowest angular momentum case A1 ($l=1.50$), the emission remains very low and nearly constant throughout the simulation. This is consistent with the absence of strong compression and variability inside the CENBOL region as found in Fig.  3 and Fig.  4. For case A2 ($l=1.60$), the luminosity rises slightly and shows small amplitude variability compared to A1. For higher $l$ cases (runs A3, A4, and A5), we observe the bolometric luminosity to increase with $l$. At the same time, all these cases show significant variability; the amplitudes vary by a few factors. Average size of the CENBOL increases with increasing $l$. Thus, increasingly larger volumes with denser matter enhance the bremsstrahlung emission. The variability is caused by the dynamical CENBOL. Fig. 4 already indicates the variability of CENBOL, and observed variability in luminosity is consistent with this.
We also investigate the spectral properties of the disk due to this bremsstrahlung emission and its self-Comptonization effect. We model the bremsstrahlung emission and (inverse-)Comptonization of the emitted photons using Monte-Carlo techniques as described above in Section 2.3. Fig. 9 shows the final spectra for all the cases. The plots show the specific intensity (in arbitrary units) as a function of photon energy (measured in keV). Since we are performing single-component fluid dynamics (i.e., electrons or protons/ions are not separated), we have a single-component temperature. Fig. 3 shows the radial profile of this single-component temperature along the equator. The temperature profiles of the fluid result in the bremsstrahlung spectra to have an exponential cut-off at very high energy, and hence, the specific intensities peak close to 5000 keV. We have chosen a low energy cut off at 0.01 keV in the spectra because the specific intensity falls off by an order of 5 below $0.01$ keV. Because of the lower density, most of the photons escape without being scattered. Among all the cases, we find the highest scattering events for case $A5$ around 0.250.25% of the seed photons got scattered. Highest optical depth is found to be 0.15 cm-1 for this case integrated along the equatorial plane(edge-on). For all the cases, the highest energy photons originate from the innermost high-temperature region. As these photons travel away from the black hole, they mostly get redshifted. Also, we find that if such high-energy photons get scattered at the outer, cooler region, they mostly deposit their energy to the electrons via Comptonization.
In passing, we would like to mention that even though our disk-jet systems produce photons of energy $\sim$ MeV, we do not consider the electron-positron pair production. The possibility of pair production in high-energy astrophysical sources is not determined solely by photon energy. Instead, it depends critically on the radiation density, which is quantified by a dimensionless parameter (Guilbert et al. 1983):
$\ell=\frac{L\sigma_{T}}{Dm_{e}c^{3}},$ (28)
where $L$ is the luminosity measured in units of Eddington luminosity, $D$ is the characteristic size (measured in $r_{g}$ unit) of the emitting region, $\sigma_{T}$ is the Thomson cross-section, and $m_{e}$ is the electron mass. For our case, the Eddington luminosity is
$L_{\rm Edd}=1.3\times 10^{38}\left(\frac{M_{bh}}{M_{\odot}}\right)\simeq 8.45\times 10^{47}\ \mathrm{erg\ s^{-1}}.$ (29)
The maximum bolometric luminosity we observe from the result $L_{\rm bol}\sim 10^{42}\ \mathrm{erg\ s^{-1}}$. Hence, $\frac{L_{\rm bol}}{L_{\rm Edd}}\sim 10^{-6}$, if we consider $D\sim 2651r_{g}$, we find $\ell\sim 10^{-9}.$ Since $\ell\ll 1$ (Svensson 1984), the photon–photon optical depth is extremely small. Although the bremsstrahlung spectrum extends up to $\sim 100$ MeV, the radiation field is too dilute for significant pair production. High-energy photons escape without generating a pair cascade.
4 Summary and Conclusions
In this paper, we study unmagnetized accretion disk-outflow system around a super-massive black hole. These simulations results are equally applicable for stellar-mass black holes, also provided the radiative efficiency is low. We simulate the accretion of nearly free-falling sub-Keplerian matter which develops a centrifugal barrier due to non-zero angular momentum. A part of the infalling matter is compressed due to this barrier and leaves the system as outflow. In this work, we investigate this mechanism and quantify the outflow by varying the angular momentum of the infalling matter. We use radiative-hydrodynamics simulations for our study, where radiative cooling due to bremsstrahlung emission is included. Our main findings and conclusions are as follows:
  • Nearly free-falling sub-Keplerian matter slows down close to a black hole due to a strong centrifugal barrier. The strength of the barrier depends upon the specific angular momentum of the flow. Because of the barrier, a shock develops in the flow and shock compressed flow forms a hotter, denser corona surrounding the black hole.
  • The shock-compressed flow subsequently accelerates towards the black hole and further gets geometrically compressed. Part of this compressed, rotating flow experiences a net outward force due to the combined effects of pressure gradient and centrifugal forces, which accelerates flow away from the black hole.
  • This accelerated matter becomes unbounded and escapes the accretion disk as a collimated, bipolar outflow out to thousands of radii. In the absence of further acceleration, this outflowing matter achieves a terminal speed that depends on the strength of the centrifugal barrier and the resulting compression. For our simulated cases, we find a maximum terminal speed of $0.14c$. Net mass flux of outflow also depends on the specific angular momentum.
  • The post-shock flow shows dynamical behavior that reflects upon the outflow and outgoing radiative properties from the accretion disk. The mass flux in the bipolar outflow and the total luminosity show significant temporal variabilities consistent with the variability seen in the CENBOL boundary at the equator.
  • The bremsstrahlung spectra are primarily determined by the innermost, dense region of the accretion flow. This makes the resulting spectra peak at high energies and produce hard spectral states. Low optical depth of the accretion disk, resulting due to low accretion rate, doesn’t show a significant Comptonization effect.
Acknowledgements.
We acknowledge the usage of the Kepler Computing facility, maintained by the Department of Physical Sciences, IISER Kolkata. We also acknowledge the usage of IUCAA’s Pegasus Computing facility for conducting a few simulations. Authors also acknowledge the help of Mr. Bungkiu Kissinquinker during the initial phase of this work.